OpenStudy (anonymous):

44 = 14 − 2a. y – 5 = –8 –7b = 21 −5 = c / 7 these are all the ones i need help with

OpenStudy (ilovecake):

44=14-2a Simple and best practice solution for 44=14-2a equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 44=14-2a equation: Simplifying 44 = 14 + -2a Solving 44 = 14 + -2a Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '2a' to each side of the equation. 44 + 2a = 14 + -2a + 2a Combine like terms: -2a + 2a = 0 44 + 2a = 14 + 0 44 + 2a = 14 Add '-44' to each side of the equation. 44 + -44 + 2a = 14 + -44 Combine like terms: 44 + -44 = 0 0 + 2a = 14 + -44 2a = 14 + -44 Combine like terms: 14 + -44 = -30 2a = -30 Divide each side by '2'. a = -15 Simplifying a = -15

OpenStudy (anonymous):

You are so patient. :) @Ilovecake

OpenStudy (ilovecake):

:) @tinybookworm